OpenStudy (anonymous):

For the questions below: A statistics student (Lee) wants to compare his final exam score to his friend Zoey’s final exam score from last year. However, the two exams were scored on different scales. Lee asks Zoey for the mean and standard deviation of her class on the exam as well as her final exam score. In relation to the other people in their classes, who had the better exam score? Lee: Final exam score = 85; Class: M = 70; SD = 10 Zoey: Final exam score = 45; Class: M = 35; SD = 5

ganeshie8 (ganeshie8):

Lee : 85 = 70 + 1.5*10 = 70 + 1.5(SD) Zoey : 45 = 35 + 2*5 = 35 + 2(SD)

ganeshie8 (ganeshie8):

that means Lee just did 1.5 SD better than mean, where as Zoey did 2 SD better than his mean

ganeshie8 (ganeshie8):

So clearly, Zoey had a better exam score